turf's grass seed cost 60 cents a pound and lumpy's grass seed cost cost 45 cents a pound. How many pounds of each were used to to make a 40-pound mixture that cost $20.25


i'm so confused!!

Answer by tutorcecilia(2144)

Use the following formulas:
x+y= total amount
ax+by= total cost
.
Plugging in the values:
x+y=40
or y = 40-x
and
.60x+.45y=20.25
.
Substituting y=40-x gives:
.60x + .45(40-x) = 20.25
.
Solving for the x-value:
.60x+18-.45x = 20.25
.60x-.45x=20.25-18
.15x =2.25
x=2.25/.15
x=15lbs at .60 cents/lbs
.
So,
x+y=40
15+y=40
y=40-15
y=25 lbs at .45 cents/lbs